# Two-dimensional Brownian motion with dependent components: turning angle analysis

### v2

@@ V Time-varying correlation
In many real-world data, additional complexities can cause the correlation coefficient between components to change over time, that is, $\rho$ may not be constant. In the financial dataset analyzed in the previous section, we limited ourselves to a 2-year period to avoid such a situation. However, data for both indices have been available for more than 67 years. Thus, in a general situation, one might find that the correlation coefficient depends on time, where $\rho(t)$ describes the local correlation between coordinates. We can express the correlation through the time-dependent cross-covariance function
@@ V.1 Detecting non-constant correlation
An important question is how to detect fluctuations in the correlation $\rho$ from experimental data. A way to achieve this goal involves the analysis of the turning angle distribution and how it changes over time. In cases where only a single trajectory is available, e.g., financial datasets, we can infer the correlation coefficient ${\rho}$ or the distribution of turning angles by employing a sliding window of length $w$. To present this approach, we consider simulated data with $\rho(t)$ corresponding to the stepwise constant function presented in Figure 8(a). In our numerical simulations, we use a moving window of length $w=400$, while the length of the sample trajectory is $2^{11}$. The results obtained are presented in Figure 9, where panel (a) shows the evolution of the histogram of the turning angles over time. The darker colors of the histograms correspond to earlier parts of the trajectory, while time $t$ on the vertical axis corresponds to the index of the midpoint of the sliding window, i.e. the histogram at $t=600$ corresponds to the turning angles calculated at times $t=401,402,\ldots,800$. Figure 9(b) shows the correlation $\widehat{\rho}(t)$ estimated in two different ways. First, the standard Pearson estimator of the correlation coefficient $\rho$ is obtained as a function of time $t$ corresponding to the data points in the window $\{t-w/2+1,\ldots,t+w/2\}$, in a similar way as in Figure 9(a). Second, $\widehat{\rho}(t)$ is calculated from the distribution of turning angles within the same window, using a least-squares method (denoted in Figure 9 as "angles"). Here, the distribution of the turning angle is fitted to the function presented in eq.  (17). Both estimation methods give a similar result, comparable to the ground truth $\rho(t)$.
@@ V.2 Example: Financial data
VI Conclusion
In this article, we have introduced and analyzed a model for two-dimensional Brownian motion with spatial correlations, characterized by the correlation coefficient $\rho$ and the diffusivities in each direction $D_{1}$ and $D_{2}$. We derived some of the key properties, such as the autocovariance and cross-covariance functions, the mean square displacement, and the turning angle distribution given in a general case by a convolution. We demonstrated that while the MSD exhibits the expected linear behavior typical of normal diffusion, spatial correlation introduces significant changes in the distribution of turning angles and the overall trajectory patterns.
Through numerical simulations, we illustrate the impact of correlation coefficients on sample trajectories, PDF contour plots, and displacement histograms. Numerical simulations confirmed that spatial correlations markedly affect the statistical properties of motion, providing a deeper insight into the behavior of the introduced model. Historically, turning angles have been underutilized outside of random walks with temporal correlations, leading to a gap in the comprehensive understanding of the data. By incorporating turning angles alongside correlations, we gain a better understanding of the data and improve our understanding with directional information. To provide an additional focus on the usefulness of the model, we applied it to real-world datasets, namely financial indices (Dow Jones Industrial Average index and Standard and Poor’s 500) and physical systems (polystyrene beads in aqueous solution). The results highlighted the model’s applicability across fields, reinforcing its utility in describing complex systems with spatial correlations.
@@ VI Conclusion
Future work could incorporate a memory structure for each of the coordinates, e.g. by using a two-dimensional fractional Brownian motion model [57, 58, 39, 59]. The analytical results for the turning angles would be more complicated to obtain, as such a model incorporates temporal correlations. However, the extension to higher dimensions of correlated BM is straightforward. The turning angles are still random variables with PDF as given in eq. (10), although dependent on a larger number of parameters. Other possible extensions relate to processes with an infinite second moment, such as Lévy flights [60, 61] and fractional Lévy stable motion [62, 63], and to processes with non-stationary increments, such as multifractional Brownian motions [64, 65, 66, 67], or processes with varying in time parameters [68], and continuous time random walks [69, 70]. Similarly, other dependence measures, such as covariation or codifference, play a similar role to the correlation coefficient in this framework [71].

### v4

@@ V Time-varying correlation
In many real-world data, additional complexities can cause the diffusion process to change over time due to temporal or spatial heterogeneities [68, 69, 70, 71, 72, 73]. In the same manner, the correlation coefficient between components can fluctuate, that is, $\rho$ may not be constant. In the financial dataset analyzed in the previous section, we limited ourselves to a 2-year period to avoid such a situation. However, data for both indices have been available for more than 67 years. Thus, in a general situation, one might find that the correlation coefficient depends on time, where $\rho(t)$ describes the local correlation between coordinates. We can express the correlation through the time-dependent cross-covariance function
@@ V.1 Detecting non-constant correlation
An important question is how to detect fluctuations in the correlation $\rho$ from experimental data. A way to achieve this goal involves the analysis of the turning angle distribution and how it changes over time. In cases where only a single trajectory is available, e.g., financial datasets, we can infer the correlation coefficient ${\rho}$ or the distribution of turning angles by employing a sliding window of length $w$. To present this approach, we consider simulated data with $\rho(t)$ corresponding to the stepwise constant function presented in Figure 9(a). In our numerical simulations, we use a moving window of length $w=400$, while the length of the sample trajectory is $2^{11}$. The results obtained are presented in Figure 10, where panel (a) shows the evolution of the histogram of the turning angles over time. The darker colors of the histograms correspond to earlier parts of the trajectory, while time $t$ on the vertical axis corresponds to the index of the midpoint of the sliding window, i.e. the histogram at $t=600$ corresponds to the turning angles calculated at times $t=401,402,\ldots,800$. Figure 10(b) shows the correlation $\widehat{\rho}(t)$ estimated in two different ways. First, the standard Pearson estimator of the correlation coefficient $\rho$ is obtained as a function of time $t$ corresponding to the data points in the window $\{t-w/2+1,\ldots,t+w/2\}$, in a similar way as in Figure 10(a). Second, $\widehat{\rho}(t)$ is calculated from the distribution of turning angles within the same window, using a least-squares method (denoted in Figure 10 as "angles"). Here, the distribution of the turning angle is fitted to the function presented in eq.  (18). Both estimation methods give a similar result. For $\rho=0.8$, the estimation of correlation coefficient from the turning angle distribution performs very well, yielding accurate results. However, as the correlation decreases ($\rho=0.5$) the estimation error increases significantly. In such cases, using a larger time window is necessary to achieve more accurate correlation estimates. A larger window would be needed for accurate correlation estimation. In any case, it is clear from the data that the distribution of turning angles, using a small time window, undergoes a noticeable change around $t=1000$, accompanied by a decrease in correlation.
@@ V.2 Example: Financial data
VI Conclusions
In this article, we have introduced and analyzed a model for two-dimensional Brownian motion with spatial correlations, characterized by the correlation coefficient $\rho$ and the diffusivities in each direction $D_{1}$ and $D_{2}$. We derived some of the key properties, such as the autocovariance and cross-covariance functions, the mean square displacement, and the turning angle distribution given in a general case by a convolution. We demonstrated that while the MSD exhibits the expected linear behavior typical of normal diffusion, spatial correlation introduces significant changes in the distribution of turning angles and the overall trajectory patterns. The distribution of turning angles offers valuable insights into the correlation between trajectory components. While such correlations can also be analyzed in terms of the Pearson correlation coefficient of the increments projected onto different axes, turning angles serve as a crucial measure in the study of dynamics within complex systems. In particular, while Pearson’s coefficient is sensitive to an arbitrary change in orientation of the coordinate system, the turning angle statistics are not. Thus, the turning angles probe the internal dynamics of the physical system, independent of the orientation of the coordinate system of the laboratory frame. These statistics are also useful for evaluating temporal correlations, which may arise in contexts such as viscoelastic environments or fractal geometries. Consequently, it is important to understand how spatial correlations (as studied in this work) influence the distribution of turning angles, as well as their impact on other metrics like the mean squared displacement (MSD) and covariance functions, and how it changes over time in heterogeneous processes.
Through numerical simulations, we illustrate the impact of correlation coefficients on sample trajectories, PDF contour plots, and displacement histograms. Numerical simulations confirmed that spatial correlations markedly affect the statistical properties of motion, providing a deeper insight into the behavior of the introduced model. Historically, turning angles have been underutilized, leading to a gap in the comprehensive understanding of the data. The turning angle distribution is slowly gaining more attention and has been employed to study a diverse array of systems including the compartmentalization and clustering in the plasma membrane [62, 64], the motion of telomeres in the nucleus [66], the dynamics of insulin granules in the cytoplasm and contractile cytoskeletal networks on model membranes [60], turbulent flows [61], and even the trajectories of football players [63]. By incorporating turning angles alongside correlations, we gain a better understanding of the data and improve our understanding of directional information. To provide an additional focus on the usefulness of the model, we applied it to real-world datasets, namely financial indices (Dow Jones Industrial Average index and Standard and Poor’s 500) and physical systems (polystyrene beads in aqueous solution). The results highlighted the model’s applicability across fields, reinforcing its utility in describing complex systems with spatial correlations.
@@ VI Conclusions
Future work could incorporate a memory structure for each of the coordinates, e.g. by using a two-dimensional fractional Brownian motion model [74, 75, 46, 76]. The analytical results for the turning angles would be more complicated to obtain, as such a model incorporates temporal correlations. However, the extension to higher dimensions of correlated BM is straightforward. The turning angles are still random variables with PDF as given in eq. (10), although dependent on a larger number of parameters. Other possible extensions relate to processes with an infinite second moment, such as Lévy flights [77, 78] and fractional Lévy stable motion [79, 80], and to processes with non-stationary increments, such as multifractional Brownian motions [81, 82, 83, 84], heterogeneous diffusion processes with transitions points [68, 69, 70, 71], and continuous time random walks [85, 86]. Similarly, other dependence measures, such as covariation or codifference, play a similar role to the correlation coefficient in this framework [87].